# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v3

@@ 4.2 Optimization
Based on the above analyses, we now introduce the training of TAD in detail. Following previous work Hafner et al. (2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
To balance different tasks, in the stage of collecting data, TAD utilizes different replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively. Then in the training stage, TAD will sample data from each replay buffer and train the reward-informed world models via the objective, which follows Eq. (3), as
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+\sum_{t=1}^{T}\ln p_{\theta}(\mathcal{M}_{i}|h_{t}^{i},s_{t}^{i})\right].\end{split}$ (4)
In Eq. (4), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. The novel last term is calculated in Eq. (3) for encouraging the agent to encode task information into hidden states and infer the current task based on these hidden states, which benefits generalizing to unseen tasks (We demonstrate this in Sec. 4.3).
In terms of training the actor-critic, both of which are parameterized neural networks, we follow the actor-critic learning approach in Dreamer with a more powerful hypothesis set $\Pi_{3}$ (The necessity is shown in Sec. 4.3). When optimizing the policy, TAD samples a series of states from the replay buffer and starts from them to imagine trajectories via interacting with reward-informed world models, which is fixed at this time. Here reward-informed world models can capture invariant features, which are beneficial for the agent to train more effectively and gain better generalization. After obtaining imagined trajectories, the actor-critic can be optimized via maximizing the $\lambda$-return Schulman et al. (2015) and regressing the targets calculated by TD Sutton and Barto (2018), respectively.
@@
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ testing tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
1:  while not converge do
@@
5:        Update the world models via optimizing Eq. (4).
@@ 4.3 Theoretical Analyses
In this part, we will provide some analyses to show that components of TAD are simple but effective for task generalization.
Is it necessary to use policies that utilize all historical information for task generalization?
Below we will introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling the task generalization.
@@
    Markovian policy set $\Pi_{1}$ Sutton and Barto (2018); Yarats et al. (2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
@@
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ Hafner et al. (2019a, 2020); Lee et al. (2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
@@
    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ Rakelly et al. (2019); Zintgraf et al. (2019), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1(a), naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. Next, we will analyze their expressive ability under task distribution $\mathcal{T}$. Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first main result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal, summarized as:
Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix B.3).
@@
To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric named Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
@@
[…]
@@
Theorem 3 (Proof in Appendix B.4).
@@
[…]
Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$. When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited. This conclusion is further verified empirically in the experimental section. Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix B.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability. Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
@@
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. Notice that $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be regard a deformation of $p(\mathcal{M}|l),l\in\mathcal{L}$ as $h_{t}$ encodes all historical information.
Theorem 4 (Informally, detailed analyses and proof are in Appendix B.6).
@@
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)dl\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\qquad-\left.\max_{a}\int p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right],\end{split}$ (7)
@@
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq. (7), thus is effective for improving the generalization ability of $\pi$. More details and discussion are also provided in Appendix B.6.
Tasks Hypothesis Walker-stand&\&&walk Acrobot-swingup&\&&sparse Reacher-easy&\&&hard Walker-walk&\&&prostrate
CURL $\Pi_{1}$ 563.4 $\pm$ 16.0 11.0 $\pm$ 1.8 552.5 $\pm$ 88.9 441.0 $\pm$ 7.7
SAC+AE $\Pi_{1}$ 683.8 $\pm$ 164.3 11.7 $\pm$ 5.4 669.5 $\pm$ 99.4 445.8 $\pm$ 2.6
PlaNet $\Pi_{2}$ 990.6 $\pm$ 3.8 12.2 $\pm$ 9.8 341.0 $\pm$ 228.3 445.4 $\pm$ 5.1
Dreamer $\Pi_{2}$ 992.3 $\pm$ 1.7 159.9 $\pm$ 108.4 506.1 $\pm$ 382.9 449.7 $\pm$ 12.5
TAD $\Pi_{3}$ 993.9 $\pm$ 1.0 217.3 $\pm$ 107.1 823.3 $\pm$ 237.4 913.7 $\pm$ 11.5
@@
Tasks Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
@@
CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 137.1 $\pm$ 8.6 57.4 $\pm$ 21.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 144.0 $\pm$ 4.5 68.6 $\pm$ 3.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 99.8 $\pm$ 28.3 80.5 $\pm$ 24.2 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 92.2 $\pm$ 19.9 79.6 $\pm$ 14.8 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
TAD $\Pi_{3}$ 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 344.8 $\pm$ 2.8 132.6 $\pm$ 14.2 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
@@
Tasks Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
Train&\&&Test Train Test Train Test

### v4

@@ 4.2 Optimization
Based on the above analyses, we now introduce the training of TAD in detail. Following previous world models (Hafner et al. 2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
To balance different tasks, when collecting data, TAD utilizes $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively. Then, TAD samples data from each replay buffer and trains the reward-informed world models via the optimization objective following Eq. (LABEL:eq_elbo) as
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (4)
In Eq. (LABEL:eq_loss), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. Besides them, TAD includes an additional task term $L_{\text{task}}$ for predicting different tasks and learning task-aware embedding. In detail, we provide two alternatives: cross-entropy and self-contrastive, which are introduced below (We theoretically demonstrate their effectiveness in Sec. 4.3).
Cross-Entropy.
The last term in Eq. LABEL:eq_elbo indicates that we need to maximize the log probability of the task context to distinguish different tasks via historical information. Thus TAD-CE directly maximizes the log probability over different tasks represented by one-hot vectors following previous works (Yang et al. 2020) and $L_{\text{task}}$ is set as below.
$\begin{split}L_{\text{task}}=\sum_{t=1}^{T}\ln p_{\theta}(m^{i}|h_{t}^{i},s_{t}^{i}).\end{split}$ (5)
Supervised-Contrastive.
Besides directly maximizing the log probability, we further propose TAD-SC, borrowing the idea of supervised contractive learning (Khosla et al. 2020). In detail, TAD-SC keeps the task embeddings of the same task closer and the task embeddings of different tasks far apart, which benefits reward-informed world models to better distinguish different tasks. Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and set $L_{\text{task}}$ as
$\begin{split}L_{\text{task}}=\sum_{j=1}^{M\times T}\sum_{a\in A(j)}\ln\frac{\exp(m_{j}\cdot m_{a}/\tau)}{\sum_{b\neq j}\exp(m_{j}\cdot m_{b}/\tau)},\end{split}$ (6)
here $A(j)$ is the set of indices that are sampled with the same task of $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works (Khosla et al. 2020).
In terms of training the actor-critic that are parameterized neural networks, we extend the actor-critic learning in Dreamer to the task distribution. We first sample a series of states from the replay buffer and start from them to imagine trajectories via our reward-informed world models (results are in Fig. 3), which can capture invariant features and are beneficial for the agent to gain better generalization. After obtaining imagined trajectories, the actor-critic is optimized via maximizing the $\lambda$-return (Schulman et al. 2015) and regressing the TD targets (Sutton and Barto 2018), respectively.
@@
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
[…]
@@
5:        Update the world models via optimizing Eq. (LABEL:eq_loss).
@@ 4.3 Theoretical Analyses
Below, we provide theoretical analyses to show that TAD’s components are simple but effective for task generalization.
Are policies that utilize all historical information in TAD necessary for task generalization?
Below we introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
@@
    Markovian policy set $\Pi_{1}$ (Sutton and Barto 2018; Yarats et al. 2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
@@
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ (Hafner et al. 2019a, 2020; Lee et al. 2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
@@
    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ (Zintgraf et al. 2019; Rimon et al. 2024), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1, naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. Now we will analyze their expressive ability under task distribution $\mathcal{T}$. Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
@@
To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
@@
$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]\\ &-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (7)
@@
Theorem 3 (Proof in Appendix A.4).
@@
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)]\end{split}$ (8)
Thus $J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{2}=J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{1}\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right]$, here $\pi^{*}=\arg\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$.
Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$. When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited. This conclusion is further verified empirically in the experimental section. Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix A.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability. Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
@@
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be rewritten as $p(\mathcal{M}|l),l\in\mathcal{L}$.
Theorem 4 (Informally, detailed analyses and proof are in Appendix A.6).
@@
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (9)
@@
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq. (LABEL:eq_pi3), thus is effective for improving the generalization ability of $\pi$. More details and discussion are also provided in Appendix A.6.
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
PlaNet $\Pi_{2}$ 309.5 $\pm$ 59.9 606.7 $\pm$ 152.9 244.8 $\pm$ 17.8
Dreamer $\Pi_{2}$ 974.2 $\pm$ 5.8 722.2 $\pm$ 12.6 241.1 $\pm$ 19.5
MAMBA $\Pi_{3}$ 994.7 $\pm$ 3.2 436.8 $\pm$ 116.1 375.6 $\pm$ 44.0
TAD-CE (Ours) $\Pi_{3}$ 998.9 $\pm$ 0.4 778.9 $\pm$ 63.1 549.6 $\pm$ 28.6
TAD-SC (Ours) $\Pi_{3}$ 982.6 $\pm$ 2.0 807.8 $\pm$ 85.2 588.8 $\pm$ 20.2
@@
Algorithms Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
@@
CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 140.2 $\pm$ 1.7 46.1 $\pm$ 29.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 130.6 $\pm$ 12.2 89.0 $\pm$ 25.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 92.5 $\pm$ 31.3 70.6 $\pm$ 18.4 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 87.8 $\pm$ 16.1 87.3 $\pm$ 20.5 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
MAMBA $\Pi_{3}$ 568.0 $\pm$ 229.1 475.7 $\pm$ 316.6 153.8 $\pm$ 34.7 121.1 $\pm$ 29.2 104.5 $\pm$ 25.7 99.9 $\pm$ 43.1
TAD-CE (Ours) $\Pi_{3}$ 937.4 $\pm$ 9.8 909.8 $\pm$ 21.9 283.9 $\pm$ 16.2 163.8 $\pm$ 53.0 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
TAD-SC (Ours) $\Pi_{3}$ 919.3 $\pm$ 21.9 906.7 $\pm$ 21.7 204.4 $\pm$ 62.2 143.9 $\pm$ 70.7 159.0 $\pm$ 36.9 104.3 $\pm$ 43.7
@@
Algorithms Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
Train&Test Train Test Train Test